Given the reaction

Cu + 4 HNO3 → Cu(NO3)2 + 2 H2O + 2 NO2


What is the total mass of NO2 produced when 26 grams of Cu is completely consumed?
1
Expert's answer
2020-02-28T12:14:01-0500

1. Find the amount of copper:

26/64=0.41 mol, v=m/M

2. Find the amount of substance NO2

0.41/1=x / 2, x=0.82

3.Find the mass NO2:

0.82*46=37.38 g